OpenStudy (anonymous):

In triangle ABC, ∠A = 80°. The bisectors of the base angles, ∠B and ∠C, intersect at point P. ∠BPC = 50° 100° 130°

OpenStudy (campbell_st):

draw a diagram  well you know B + C = 180 - 80 so B + C = 100 if both B and C are bisected then the measure of 1/2 (B + C) = 50 so using angle sum of a triangle BPC = 180 - 1/2 (B + C) I'll leave you to get the final answer

OpenStudy (anonymous):

thanks